On a day when the temperature reaches 50c, what is the temperature in degrees F and in Kelvins

1
Expert's answer
2021-11-16T11:24:48-0500

F= C*1,8 +32=50*1,8+32=122 F

Kelvins= C +273,15=50+273,15=323,15 K